Convert the following temperatures to the Celsius scale.
   293 K

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

293 K into °C
293 – 273 = 20°C

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
To convert K into °C , we need to subtract 273.
293 K into °C
= 293 - 273 = 20°C
 
Hope it helps!
Thank you ✌️